# Optimal Estimation of Temperature

### v4

@@ Appendix E The approximate expression of $\hat{\beta}_{B}$ and $\hat{T}_{G}$ in large-$N$ limit
$\begin{split}\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-1}\\ \text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=-\frac{\partial^{2}\left\langle E\right\rangle}{\partial\beta^{2}}\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-3}\\ \text{d}^{3}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-5}\left[3\left(\frac{\partial^{2}\left\langle E\right\rangle}{\partial\beta^{2}}\right)^{2}-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\frac{\partial^{3}\left\langle E\right\rangle}{\partial\beta^{3}}\right]\end{split}$ (55)
Then, substituting Eq. (55) into Eq. (54), we obtain the variance of $\hat{\beta}_{B}$,
$\Delta\hat{\beta}_{B}^{2}=\frac{1}{(\Delta E)^{2}}+\frac{1}{2}\frac{\left\langle(\delta E)^{3}\right\rangle^{2}}{(\Delta E)^{8}}+O(N^{-3}).$ (56)
@@
$S_{G}(E)=S_{B}(E)+\ln\hat{T}_{G}(E)+O(N^{-1}),$ (57)
Eq. (57) is consistent with the exact relation $\hat{T}_{G}=\hat{T}_{B}(1-C_{G}^{-1})$ , where $C_{G}=\partial E/\partial\hat{T}_{G}$ [1]. Taking the derivative of Eq. (57) with respect to $E$ on both sides and using Eq. (52), we have
$\hat{T}_{G}(E)=\hat{T}_{L}(E)-\frac{\hat{T}_{L}(E)^{2}}{2}\frac{\text{d}^{2}_{E}\hat{T}_{L}(E)}{\text{d}_{E}\hat{T}_{L}(E)}+O(N^{-2}),$ (58)
Following the similar procedure, we obtain the variance of $\hat{T}_{G}$ using Eqs. (55) (58) and $\hat{T}_{L}\left(\left\langle E\right\rangle\right)=T$,
$\Delta T^{2}_{G}=\frac{T^{4}}{(\Delta E)^{2}}+\frac{T^{12}(\text{d}^{2}_{T}\left\langle E\right\rangle)^{2}}{2(\Delta E)^{8}}+O(N^{-3}).$ (59)
@@ Appendix F Estimation of temperature for the classical ideal gas
$Z=\int_{0}^{\infty}\sigma(E)\text{e}^{-\beta E}=\beta^{-\frac{3N}{2}}\Gamma\left(\frac{3N}{2}\right),$ (60)
@@
$E^{*}=\frac{3N-2}{2\beta},$ (61)
the Boltzmann temperature
$\hat{\beta}_{B}=\frac{3N-2}{2E},$ (62)
and the Gibbs temperature
$\hat{T}_{G}=\frac{2E}{3N}.$ (63)
Using Eq. (60), we obtain the average energy
$\left\langle E\right\rangle=-\frac{\partial\ln Z}{\partial\beta}=\frac{3N}{2\beta}.$ (64)
@@
$C=\frac{\partial\left\langle E\right\rangle}{\partial T}=\frac{3N}{2}$ (65)
@@
$\hat{T}_{L}(E)=\frac{2E}{3N},\quad\hat{\beta}_{L}(E)=\frac{3N}{2E}.$ (66)
@@
$\begin{split}&\left(\frac{3N}{2}-1\right)\left(\ln\frac{3N}{2}-\ln\beta\right)\\ =&\left(1-\frac{3N}{2}\right)\ln\beta+\ln\Gamma\left(\frac{3N}{2}\right)-\frac{1}{2}\ln 2\pi\\ &-\frac{1}{2}\ln\left(\frac{3N}{2}\right)+\left(\frac{3N}{2}\right)+O(N^{-1}),\end{split}$ (67)
where we have used the String approximation [87] in the derivation.
Moreover, the left-hand side (LHS) and the right-hand side (RHS) of Eqs. (12) and (15) are exactly equal, because our model is sufficiently simple. In general, they are approximately valid.
@@
$P(E)=\frac{\beta^{\frac{3N}{2}}E^{\frac{3N}{2}-1}\text{e}^{-\beta E}}{\Gamma\left(\frac{3N}{2}\right)}.$ (68)
Using Eq. (68), we obtain the standard deviation of Boltzmann temperature
$\Delta\hat{\beta}_{B}=\beta\sqrt{\frac{2}{3N-4}},$ (69)
@@
$\Delta E=\beta^{-1}\sqrt{\frac{3N}{2}},$ (70)
@@
$\mu_{3}=\sqrt{\frac{8}{3N}}.$ (71)
Thus, the result of the achievable ETU (Eq. (13)) using Eqs. (69) (70) is:
$\begin{split}\Delta\hat{\beta}\Delta E&\geq\Delta\hat{\beta}_{B}\Delta E=\sqrt{\frac{3N}{3N-4}}>1.\end{split}$ (72)
@@
$\begin{split}\sqrt{\frac{3N}{3N-4}}&=1+\frac{2}{3N}+O(N^{-2}).\end{split}$ (73)
@@
$\Delta\hat{T}_{G}=\beta^{-1}\sqrt{\frac{2}{3N}},\quad\Delta E=\beta^{-1}\sqrt{\frac{3N}{2}},$ (74)
@@
$C=\frac{\partial\left\langle E\right\rangle}{\partial T}=\frac{3N}{2}.$ (75)

### v5

@@ Appendix E The approximate expression of $\hat{\beta}_{B}$ and $\hat{T}_{G}$ in large-$N$ limit
$\begin{split}\text{d}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-1}\\ \text{d}^{2}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=-\frac{\partial^{2}\left\langle E\right\rangle}{\partial\beta^{2}}\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-3}\\ \text{d}^{3}_{\left\langle E\right\rangle}\hat{\beta}_{L}&=\left(\frac{\partial\left\langle E\right\rangle}{\partial\beta}\right)^{-5}\left[3\left(\frac{\partial^{2}\left\langle E\right\rangle}{\partial\beta^{2}}\right)^{2}-\frac{\partial\left\langle E\right\rangle}{\partial\beta}\frac{\partial^{3}\left\langle E\right\rangle}{\partial\beta^{3}}\right].\end{split}$ (E8)
For convenience of calculation, we list the following expressions: $\partial\left\langle E\right\rangle/\partial\beta=-\left\langle(\delta E)^{2}\right\rangle$, $\partial^{2}\left\langle E\right\rangle/\partial\beta^{2}=\left\langle(\delta E)^{3}\right\rangle$ and $\partial^{3}\left\langle E\right\rangle/\partial\beta^{3}=3(\left\langle(\delta E)^{2}\right\rangle)^{2}-\left\langle(\delta E)^{4}\right\rangle$. Then, substituting Eq. (E8) into Eq. (E7), we obtain the variance of $\hat{\beta}_{B}$,
$\Delta\hat{\beta}_{B}^{2}=\frac{1}{\left\langle(\delta E)^{2}\right\rangle}+\frac{1}{2}\frac{\left\langle(\delta E)^{3}\right\rangle^{2}}{(\left\langle(\delta E)^{2}\right\rangle)^{4}}+O(N^{-3}).$ (E9)
@@
$S_{G}(E)=S_{B}(E)+\ln\hat{T}_{G}(E)+O(N^{-1}),$ (E10)
Eq. (E10) is consistent with the exact relation $\hat{T}_{G}=\hat{T}_{B}(1-C_{G}^{-1})$ , where $C_{G}=\partial E/\partial\hat{T}_{G}$ [11]. Taking the derivative of Eq. (E10) with respect to $E$ on both sides and using Eq. (E5), we have
$\hat{T}_{G}(E)=\hat{T}_{L}(E)-\frac{\hat{T}_{L}(E)^{2}}{2}\frac{\text{d}^{2}_{E}\hat{T}_{L}(E)}{\text{d}_{E}\hat{T}_{L}(E)}+O(N^{-2}),$ (E11)
Following the similar procedure, we obtain the variance of $\hat{T}_{G}$ using Eqs. (E8) (E11) and $\hat{T}_{L}\left(\left\langle E\right\rangle\right)=T$,
$\Delta T^{2}_{G}=\frac{T^{4}}{\left\langle(\delta E)^{2}\right\rangle}+\frac{T^{12}(\text{d}^{2}_{T}\left\langle E\right\rangle)^{2}}{2(\left\langle(\delta E)^{2}\right\rangle)^{4}}+O(N^{-3}).$ (E12)
@@ Appendix F Estimation of temperature for the classical ideal gas
$Z=\int_{0}^{\infty}\sigma(E)\text{e}^{-\beta E}=\beta^{-\frac{3N}{2}}\Gamma\left(\frac{3N}{2}\right),$ (F1)
@@
$E^{*}=\frac{3N-2}{2\beta},$ (F2)
the Boltzmann temperature and the Gibbs temperature
$\hat{\beta}_{B}=\frac{3N-2}{2E},\hat{T}_{G}=\frac{2E}{3N}.$ (F3)
Using Eq. (F1), we obtain the average energy
$\left\langle E\right\rangle=-\frac{\partial\ln Z}{\partial\beta}=\frac{3N}{2\beta},$ (F4)
@@
$C=\frac{\partial\left\langle E\right\rangle}{\partial T}=\frac{3N}{2},$ (F5)
@@
$\hat{T}_{L}(E)=\frac{2E}{3N},\quad\hat{\beta}_{L}(E)=\frac{3N}{2E}.$ (F6)
@@
$\begin{split}\left(\frac{3N}{2}-1\right)\left(\ln\frac{3N}{2}-\ln\beta\right)=&\left(1-\frac{3N}{2}\right)\ln\beta+\ln\Gamma\left(\frac{3N}{2}\right)-\frac{1}{2}\ln 2\pi\\ &-\frac{1}{2}\ln\left(\frac{3N}{2}\right)+\left(\frac{3N}{2}\right)+O(N^{-1}),\end{split}$ (F7)
where we have used the String approximation [75] in the derivation. Moreover, the left-hand side (LHS) and the right-hand side (RHS) of Eqs. (12) and (15) are exactly equal, because our model is sufficiently simple.
@@
$P(E)=\frac{\beta^{\frac{3N}{2}}E^{\frac{3N}{2}-1}\text{e}^{-\beta E}}{\Gamma\left(\frac{3N}{2}\right)}.$ (F8)
Using Eq. (F8), we obtain the standard deviation of Boltzmann temperature
$\Delta\hat{\beta}_{B}=\beta\sqrt{\frac{2}{3N-4}},$ (F9)
@@
$\Delta E=\beta^{-1}\sqrt{\frac{3N}{2}},$ (F10)
@@
$\mu_{3}=\sqrt{\frac{8}{3N}}.$ (F11)
Thus, the result of the achievable ETU (Eq. (13)) using Eqs. (F9) (F10) is:
$\begin{split}\Delta\hat{\beta}\Delta E&\geq\Delta\hat{\beta}_{B}\Delta E=\sqrt{\frac{3N}{3N-4}}>1.\end{split}$ (F12)
@@
$\begin{split}\sqrt{\frac{3N}{3N-4}}&=1+\frac{2}{3N}+O(N^{-2}).\end{split}$ (F13)
@@
$\Delta\hat{T}_{G}=\beta^{-1}\sqrt{\frac{2}{3N}},\quad\Delta E=\beta^{-1}\sqrt{\frac{3N}{2}},$ (F14)
@@
$C=\frac{\partial\left\langle E\right\rangle}{\partial T}=\frac{3N}{2}.$ (F15)